nolegirl24:

Which shows 54^2 − 46^2 being evaluated using the difference of squares method? 54^2 − 46^2 = (2916 + 2116)(2916 − 2116) = 4,025,600 54^2 − 46^2 = (54 + 46)(54 − 46) = (100)(8) = 800 54^2 − 46^2 = 2916 − 2116 = 800 54^2 − 46^2 = (54 − 46)^2 = 8^2 = 64

dude:

Applying the difference of squares method \((a^2-b^2)=(a+b)(a-b)\)

nolegirl24:

so it would be like this? (54^2 - 46^2) = (54 + 46)(54 - 46)

dude:

Right

dude:

Straight off the bat leads to one of the answers

nolegirl24:

B?

dude:

Yep

dude:

56+46 = 100 54-46 =8 Since they were in parenthesis you multiply (100)(8)= 800

nolegirl24:

YAY I am done :O!!! Thanks!!!

dude:

Ofcs :P

nolegirl24:

I got a 86%!!!

nolegirl24:

I passed the review!!

dude:

o.o I got something wrong?

nolegirl24:

Nerp! You got all the ones we did right! It was the ones I did on my own that i thought i knew!!

dude:

Oh okay (was getting worried there)

dude:

Okay if you need to review I am here ig, take care!

nolegirl24:

No no you are good! Will you be on tomorrow? I have 3 to do tomorrow but they are essay questions

dude:

Yeah I have a half-school day tomorrow I should be on for most of the day

nolegirl24:

Okay! Talk to you tomorrow! :O

dude:

Take care and sleep well :P

nolegirl24:

You too ^~^